# A Generalized Difference-in-Differences Estimator for Unbiased Estimation of Desired Estimands from Staggered Adoption and Stepped-Wedge Settings

### v1

@@ 4 Data Example: Stepped Wedge Trial
2 0.903 0.721 0.808 0.411 -2.38% 0.821 -4.01% 0.540
3 0.758 0.341 0.810 0.504 -3.97% 0.496 -1.73% 0.795
4 0.683 0.012 0.736 0.044 -11.64% 0.100 -8.65% 0.080
5 0.832 0.190 0.902 0.611 -6.19% 0.147 -3.24% 0.603
6 0.716 0.062 0.750 0.182 -10.93% 0.038 -9.02% 0.173
7 0.835 0.514 0.812 0.511 -5.75% 0.634 -6.15% 0.616
@@ 5 Discussion
The proposed method provides a highly flexible and adaptable approach to analyzing stepped wedge and staggered adoption settings. In particular, it allows the investigator to specify a target estimand that is any linear combination of unique treatment effects and guarantees unbiasedness under the standard assumptions and the particular treatment effect heterogeneity assumptions specified. Among this class of unbiased estimators, the investigator can then identify the one with the lowest variance for an assumed working covariance structure. Analogous to the working correlation matrix used in generalized estimating equations (Zeger & Liang (1986)), this working covariance structure need not be correctly specified to maintain unbiased estimation. A misspecified structure, however, may lead to suboptimal variance and reduced power to detect an effect.
@@
The composition of the estimator from two-by-two DID estimators allows both intuitive connections to methods commonly used in econometrics and quantitative social science for quasi-experimental data as well as assumptions that are familiar to those stepped in that literature. In addition, it allows clear formulation of the treatment effect heterogeneity assumptions that are made rather than those assumptions arising implicitly through a linear regression formulation. By specifying the variance assumption separately, it also removes the need for a correct joint specification of the mean and variance models and allows those assumptions (for the working structure) to be formulated independently of the treatment effect assumptions. These approaches also allow for ease of sensitivity analysis, as the expectation can be found using Equation 5 using a specified weight vector under any other assumption setting. This, along with the potential to assess relative efficiency under the working covariance structure as shown in the example, allows the bias-variance tradeoff to be made explicit, and investigators can consider the appropriate generalizability and interpretation of their target estimands.
@@
In the example re-analyzing the data collected by Trajman et al. (2015), I demonstrated the use of this method for a variety of target estimands in a SWT. Results largely align with those of existing non-parametric estimators, which already had been shown to have desirable bias and inference properties (Thompson et al. (2018); Kennedy‐Shaffer et al. (2020)). In nearly comparable situations, like the assumption of a homogeneous effect compared to the inverse-variance weighted average vertical approach (Thompson et al. (2018)) or either crossover-type estimator (Kennedy‐Shaffer et al. (2020)), this approach had a lower p-value, which may indicate increased power to detect the effect. In addition, many other estimates are tractable using this approach, allowing for flexible pre-specified or exploratory analysis of treatment effect heterogeneity.
@@
Although simulation and testing in other settings are still necessary to determine overall operating characteristics of these estimators, it is reasonable that they are more powerful under homogeneity assumptions than purely vertical approaches since they use horizontal comparisons and than pure crossover approaches since they use information from other comparisons with expectation zero. In addition, the fact that the method generally highly weights cells with high information content (Kasza & Forbes (2019); Kasza, Taljaard & Forbes (2019)) and respects centrosymmetry under appropriate assumptions (Bowden et al. (2021)) indicates efficient use of the observed data. However, their performance compared to regression-based approaches (e.g., Hooper et al. (2016); de Chaisemartin & D’Haultfœuille (2020); Sun & Abraham (2021); Lindner & McConnell (2021); Kenny et al. (2022); Maleyeff et al. (2023); Lee & Cheung (2024)) and existing robust estimators (e.g., Hughes et al. (2020); Callaway & Sant’Anna (2021); Roth & Sant’Anna (2023); Borusyak et al. (2024)) remains to be determined. In particular, future study should determine when this approach provides equivalent estimators to any of those approaches (in particular, those proposed by Sun & Abraham (2021) and Lindner & McConnell (2021)) and under what settings it shows more or less desirable operating characteristics to the others.
@@
Additional future work should consider identifying other approaches to inference, including the possibility of closed-form variance estimators under certain variance assumptions. Identifying appropriate plug-in estimators for use in Equation 10, their robustness to misspecification of the working covariance structure, and the asymptotic distribution would be key, as has been done for mixed effects models with misspecification (see, e.g., Voldal et al. (2022)). This would allow for more targeted design of experiments when analysis will proceed using these methods. Understanding likely covariance structures, especially given planned or conducted approaches for sampling of units (see, e.g., Hooper (2021)), would improve selection of $\bm{M}$ as well. Additionally, modifications of this method to accommodate adjustment for covariates would be useful for non-randomized staggered adoption settings where parallel trends holds only conditionally (see, e.g., Callaway & Sant’Anna (2021); Roth et al. (2023)) and SWTs with restricted, stratified, or matched randomization (see, e.g., Copas et al. (2015)).
@@
Both stepped wedge and staggered adoption settings are common study designs with important roles to play in assessing a variety of policies and implementation approaches; for example, both have been proposed as useful tools for rapid policy evaluation in the COVID-19 pandemic or similar settings (see, e.g., Goodman-Bacon & Marcus (2020); Bell et al. (2021); Cowger et al. (2022); Kennedy-Shaffer (2024)). They each, however, have statistical intricacies that should be carefully considered in the design and analysis (see, e.g., Baker et al. (2022); Nevins et al. (2023)). The generalized DID estimator proposed here has desirable bias and variance properties and interpretability that enable its use across many such settings and desired treatment effect estimands.
@@
All code used for this manuscript is available at https://github.com/leekshaffer/GenDID. Included there is a simulated data set based on this data example. Access policies for the original data set can be found at Trajman et al. (2015), https://doi.org/10.1371/journal.pone.0123252. The author wishes to thank Prof. Anete Trajman for making the data available to the author for re-analysis and Prof. Jennifer Thompson for helpful discussions regarding the data.

### v3

@@ 5 Data Example: Stepped Wedge Trial
2 0.903 0.692 0.808 0.414 -2.38% 0.790 -4.01% 0.543
3 0.758 0.369 0.810 0.527 -3.97% 0.514 -1.73% 0.808
4 0.683 0.008 0.736 0.047 -11.64% 0.012 -8.65% 0.084
5 0.832 0.193 0.902 0.595 -6.19% 0.150 -3.24% 0.578
6 0.716 0.062 0.750 0.171 -10.93% 0.030 -9.02% 0.157
7 0.835 0.518 0.812 0.516 -5.75% 0.635 -6.15% 0.636
We can also compare these results with those from the existing methods described in Section 4. These are summarized in Table 6.
Table 6: Risk difference estimates and permutation test p-values for various staggered adoption methods and comparable generalized DID estimates for XpertMTB/RIF testing on tuberculosis outcomes in Brazil, 2012. Note: generalized DID estimators are for the most comparable estimand and assumptions, using an exchangeable correlation structure with ICC 0.003.
Comparison Method Generalized DID
Method Estimate P-Value Estimator Estimate P-Value
TW -7.35% 0.018 $\hat{\theta}^{(5)}$7.35% 0.020
CS (simple ATT) -6.41% 0.214 $\frac{1}{21}\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}_{j,a}^{(2)}$6.11% 0.229
CS (dynamic ATT) -5.39% 0.472 $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}_{a}^{(3)}$5.57% 0.456
CS (group ATT) -7.10% 0.195 $\frac{1}{6}\sum_{a=1}^{6}\sum_{j=a+1}^{7}\frac{1}{7-j+a}\hat{\theta}_{j,a}^{(2)}$8.38% 0.130
CS (calendar ATT) -5.63% 0.194 $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}_{j}^{(4)}$6.81% 0.054
SA (ATT) -10.64% 0.072 $\frac{1}{21}\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}_{j,a}^{(2)}$6.11% 0.229
CH -7.34% 0.055 $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}_{j,1}^{(2)}$8.41% 0.015
CO-1 -7.34% 0.055 $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}_{j,1}^{(2)}$8.41% 0.015
CO-2 -6.97% 0.051 $\hat{\theta}_{1}^{(3)}$6.80% 0.043
CO-3 -7.00% 0.043 $\hat{\theta}^{(5)}$7.35% 0.020
NP (time-averaged) -4.06% 0.185 - - -
NP (ATT) -4.86% 0.214 - - -
NP (inverse-variance) -4.64% 0.067 - - -
6 Data Example: Non-Randomized Staggered Adoption
In addition to using these methods in the randomized stepped-wedge trial case, we can use them to analyze observational (non-randomized) staggered adoption settings. Doing so, however, requires assessing additional assumptions and considering how the effect heterogeneity and variance assumptions comport with the assumptions necessary to interpret statistical parameters as causal parameters (i.e., Assumptions 13), especially the parallel trends assumption. The trade-offs of different approaches must be considered in this larger context for observational settings.
Several U.S. states implemented lottery-type financial incentives for COVID-19 vaccination in the spring and summer of 2021 to encourage uptake of the vaccines that had recently become available. Studies examining the effects of these policies using a variety of quasi-experimental methods generally found small positive effects, often not statistically significant for a single state (see, e.g., Sehgal (2021), Brehm et al. (2022), Fuller et al. (2022), Thirumurthy et al. (2022), Kim et al. (2023), and Lang et al. (2023)). For example, Brehm et al. (2022) identified a modest positive impact of the Ohio vaccine lottery using both DID and synthetic control methods, and Fuller et al. (2022) found that it increased uptake of the first dose of the vaccine, but not of the complete series.
I re-evaluate the impact of these policies in Midwestern states, using data collected and made publicly available by Fuller et al. (2022). I restrict the analysis to the twelve states within the Midwest geographic region delineated by the U.S. Centers for Disease Control and Prevention (National Center for Health Statistics (2024)). The outcome variable is the cumulative percentage of the adult (18+ years old) population who received at least one dose of any COVID-19 vaccine, recorded weekly for the weeks ending May 15 through July 31, 2021 (MMWR weeks 15–30, according to Morbidity and Mortality Weekly Report (2019)). These restrictions are made to both increase the plausibility of the parallel trends assumption (by restricting to more geographically similar states), reduce the number of concurrent interventions, and focus on shorter-term impacts of these policies. Per Fuller et al. (2022), four of these twelve states enacted lotteries during this time period: Ohio (MMWR Week 19), Illinois (24), Michigan (26), and Missouri (29). Note that lotteries are considered to be in effect starting in the week the policy was announced. Depending on the exact date, this may lead to a first week of intervention with fewer than seven days affected by the policy. Since the focus here is on the proposed estimators, I do not fully assess the necessary assumptions, but refer interested readers to the cited studies. Results here should be interpreted with caution.
Since the details of the lottery differ by state, as do the background conditions of the pandemic and vaccine uptake, it is reasonable to expect cluster-level heterogeneity, as well as both calendar- and exposure-time heterogeneity of treatment effects. Since calendar- and exposure-time heterogeneity imply distinct treatment effects for each treated cluster-period in this setting (because each cluster has a different adoption week), we can use assumption S2 to capture these heterogeneities. In addition, since the outcome is cumulative, an auto-regressive structure is the most reasonable, so an AR(1) correlation structure is used with first-difference parameter $\rho=0.95$, estimated from the states in other regions in those same time periods.
Table 7 displays the estimates and permutation test p-values for several summary estimands of potential interest in this setting. The weightings used for the overall ATT estimator (equal weight across all of the effects) and for the average 2–4-week effect (equal weight across the treatment effects in adoption weeks 2–4 in the three states with at least four weeks of treatment) are shown in Figure 2. Note that the staircase pattern of weights in the SWT example is not present here because the adoptions are not regularly spaced, but we can still identify which observations in the treated and untreated clusters are contributing the most.
Table 7: Cumulative percentage difference estimates and permutation test p-values for generalized DID estimators of summary effects of vaccine lotteries on COVID-19 vaccine uptake, United States Midwest region, May–July 2021. Estimation occurred using an AR(1) working correlation structure with parameter $\rho=0.95$.
Description Estimator Estimate P-Value
Overall ATT $\frac{1}{26}\sum_{\{j,a\}}\hat{\theta}_{j,a}^{(2)}$ 0.537 0.439
First-Week Effect $\frac{1}{4}\sum_{j}\hat{\theta}_{j,1}^{(2)}$ 0.285 0.155
Second-Week Effect $\frac{1}{4}\sum_{j}\hat{\theta}_{j,2}^{(2)}$ 0.605 0.065
Average Four-Week Effect $\frac{1}{12}\sum_{\{j,a:~{}j-a\leq 26,1\leq a\leq 4\}}\hat{\theta}_{j,a}^{(2)}$ 0.483 0.275
Average 2–4-Week Effect $\frac{1}{9}\sum_{\{j,a:~{}j-a\leq 26,2\leq a\leq 4\}}\hat{\theta}_{j,a}^{(2)}$ 0.561 0.276
State-Averaged Effect $\frac{1}{4}\sum_{t=15}^{30}\frac{1}{31-t}\sum_{a=1}^{31-t}\hat{\theta}_{a+t,a}^{(2)}$ 0.612 0.250
Ohio Effect $\frac{1}{12}\sum_{a=1}^{12}\hat{\theta}_{a+18,a}^{(2)}$ 0.073 0.888
Illinois Effect $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}_{a+23,a}^{(2)}$ 1.787 0.058
Figure 2: Weights on cluster-period observations for the overall ATT estimator (top) and average 2–4-week effect estimator (bottom) under assumption S2 fitted under an AR(1) working correlation structure with parameter $\rho=0.95$.
The results of the analysis comport broadly with the results found in similar lottery studies, especially Fuller et al. (2022), with modest but non-significant positive effects overall, especially in the first few weeks after full implementation. The state-specific results are also directionally consistent, with Illinois and Missouri having the largest positive estimates, Ohio a nearly zero estimated effect, and Michigan a small negative estimate. All states individually and considered together have larger estimates for the effect in the second week of the policy (denoted “Second-Week Effect” in Table 7) compared to in the first adoption week or to the overall effect. This may indicate a strong treatment effect in the first full week of the policy which then wanes and zeroes out over time, consistent with the hypothesis of some authors (e.g., Brehm et al. (2022)) that the primary effect is to pull vaccinations that would have happened anyway earlier in time. The ability to examine state- and time-heterogeneity is thus particularly important for this research question.
We can additionally compare these results to those under existing staggered adoption methods. A two-way fixed-effects analysis gives a much larger estimated treatment effect of 1.703. This is hard to interpret, however, because of the composition across the time-varying treatment effects. The CS and SA methods all give estimates between 0.45 and 0.60, with permutation p-values between 0.25 and 0.55. These are broadly similar to many of the summary effects estimated using the generalized DID. The first-period effect estimators all estimate effects around 0.22, with p-values around 0.2. Again, this is similar to the first-period effect estimated by the generalized DID, although the latter has a lower p-value, which may reflect higher power.
The working covariance matrix chosen has a larger effect on the estimates here than in the SWT case. This is likely because of the strong auto-correlation of a cumulative outcome and the large heterogeneity between clusters. See Table 13 and Figure 6 in Appendix D for estimates of the same estimands and heat maps of observation weights, respectively, under an independent working correlation structure. Note that the observation weights under AR(1) shown in Figure 2 have higher weight immediately prior to adoption, making them more similar to the CS and SA staggered adoption analysis methods. Additionally, the parallel trends assumption needs to be carefully considered. In this area, existing estimators may have advantages as they restrict comparison groups more strictly than does the generalized DID. So the investigator would need to weigh the trade-offs of potentially higher precision from incorporating more comparisons against the bias introduced by potential violations of parallel trends, preferably resulting in a pre-specified analysis plan.
7 Simulation Results
To assess the operating characteristics of these estimators, I conduct a simulation study based on the SWT described in Section 5 and the simulation study conducted in Kennedy‐Shaffer et al. (2020). An SWT with 14 clusters and 8 time periods is simulated, where two clusters cross over to the intervention at each period starting in period 2. For cluster $i$ in period $j$, the outcomes for 100 individuals are generated according to the model:
$\mu_{ij}$ $=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{ij}X_{ij},$
$Y_{ijk}$ $\overset{i.i.d.}{\sim}N(\mu_{ij},\sigma^{2}_{e}),$
where $Y_{ijk}$ is the outcome of individual $k$ within cluster $i$ in period $j$, $X_{ij}$ is an indicator of treatment for cluster $i$ in period $j$, and $\nu_{ij}\overset{i.i.d}{\sim}N(0,\sigma^{2}_{\nu})$. The following are used as fixed parameters: $\sigma^{2}_{\nu}=0.01^{2}$, $\mu=0.30$, $\bm{b}=\begin{pmatrix}0&0.08&0.18&0.29&0.30&0.27&0.20&0.13\end{pmatrix}$, and $\sigma^{2}_{e}=0.1^{2}$. The random effects for cluster, $\alpha_{i}$, are sampled without replacement from:
$\bm{a}=\begin{pmatrix}-0.016,&-0.012,&-0.011,&-0.007,&-0.005,&-0.003,&-0.001,\\ 0,&0.002,&0.003,&0.005,&0.008,&0.017,&0.020\end{pmatrix}$
These values have a mean of 0 and a standard deviation around 0.01. This generating process reflects a fixed set of clusters randomly assigned to intervention order, matching the randomization test approach used for inference.
The treatment effects $\theta_{ij}$ are varied by simulation scenario, with scenarios 1–3 having homogeneous treatment effects (akin to assumption S5), scenarios 4–6 having heterogeneous treatment effects by calendar time only (akin to assumption S4), and scenarios 7–9 having heterogeneous treatment effects by exposure time only (akin to assumption S3). Details are in Table 8.
Table 8: Treatment effect values used in the simulation scenarios. For calendar-time heterogeneous scenarios, the treatment effects are for periods starting with period 2; for exposure-time heterogeneous scenarios, the treatment effects are for the periods since intervention, starting with the first intervention period.
Scenario Heterogeneity $\theta_{ij}$ Values
1 Homogeneous $\theta^{(5)}=0$
2 Homogeneous $\theta^{(5)}=-0.02$
3 Homogeneous $\theta^{(5)}=-0.04$
4 Calendar-Time $\theta^{(4)}_{j}=\begin{pmatrix}-0.07&-0.05&-0.03&-0.01&0.01&0.03&0.05\end{pmatrix}$
5 Calendar-Time $\theta^{(4)}_{j}=\begin{pmatrix}-0.07&-0.06&-0.04&0&0.03&0.02&0.01\end{pmatrix}$
6 Calendar-Time $\theta^{(4)}_{j}=\begin{pmatrix}-0.03&-0.03&-0.03&-0.03&0&0&0\end{pmatrix}$
7 Exposure-Time $\theta^{(3)}_{a}=\begin{pmatrix}-0.010&-0.015&-0.020&-0.025&-0.030&-0.035&-0.040\end{pmatrix}$
8 Exposure-Time $\theta^{(3)}_{a}=\begin{pmatrix}0&0&-0.03&-0.03&-0.03&-0.03&-0.03\end{pmatrix}$
9 Exposure-Time $\theta^{(3)}_{a}=\begin{pmatrix}-0.07&-0.05&-0.03&-0.01&0.01&0.03&0.05\end{pmatrix}$
For each scenario, estimators using the correct assumption setting for that scenario are considered, as well as those with more flexible assumptions to determine the effect on efficiency of more robust methods. The estimators are categorized by their assumption setting and type (“GD” represents a generalized DID estimator, “CPI” an estimator using a mixed effects model with a cluster-time random effect, and “SA” an estimator from the existing staggered adoption literature). All generalized DID estimators were calculated assuming independent observations; results for an exchangeable covariance structure were very similar. Full details on the estimators used are in Appendix C. For Scenarios 1–3, all estimators are included. Scenarios 4–9 exclude estimators targeting the other heterogeneity only. All scenarios were simulated 1000 times. Mean estimates plus or minus one standard deviation are plotted in Figure 3. For each simulation, a randomization-based hypothesis test was conducted with 250 permutations; those with p-values less than $\alpha=0.05$ were considered statistically significant, and the empirical power results (or Type I Error for null effects) for each are plotted in Figure 4. Appendix D includes results for estimators targeting specific period effects.
Figure 3: Simulation results (mean estimate plus or minus one standard deviation of estimates across simulations) for 1,000 simulations each, by simulation scenario and estimator. Dotted horizontal lines represent the true value of the target estimand.
Figure 4: Simulation results (empirical power at $\alpha=0.05$ level) for 1,000 simulations each, by simulation scenario and estimator. Note that biased estimators or those targeting substantially different estimands are excluded.
In general, across simulations, most estimators were unbiased for the target estimand. When assuming full treatment homogeneity (Assumption S5), however, the estimators do not target an estimand within the convex hull of treatment effects when treatment effect varies by exposure time (so-called “dynamic treatment effects”, Scenarios 7–9). Estimators using the generalized DID method, two-way fixed effects (Goodman-Bacon (2021)), or the crossover method CO-3 (Kennedy‐Shaffer et al. (2020)) target a linear combination of treatment effects that can be identified (see Estimators 1, 2, and 4, respectively, in Figs. 3 and 4); the mixed effects model estimator appears to target a different estimand that is not currently identified. In addition, with dynamic treatment effects, a mixed effects model fully saturated with calendar time- and exposure time-heterogeneous treatment effects averaged across exposure times (Estimator 15 in Figs. 3 and 4) is not unbiased for the exposure-time-averaged treatment effect.
Aside from those cases, general patterns appear for the estimators. In general, for a certain estimand and assumption setting, existing staggered adoption methods are less efficient and less powerful than generalized DID estimators, which, in turn, are less efficient and less powerful than mixed effects model estimators. For targeting specific period effects or averages, there are cases where the generalized DID estimator can be more efficient than the corresponding mixed effects model estimator; e.g., estimators 14 and 15 in Figs. 3 and 4 and estimators 7, 8, 12, and 13 in Figs. 7 and 8. Note that for all settings considered here, the mixed effects model is (nearly) correctly specified in both its fixed effect and random effect terms. If the random effect term is misspecified to a greater extent, there may be either bias or, more commonly, reduced efficiency of these estimators.
@@ 8 Discussion
The proposed method provides a highly flexible and adaptable approach to analyzing stepped wedge and staggered adoption settings. In particular, it allows the investigator to specify a target estimand that is any linear combination of unique treatment effects and guarantees unbiasedness under the standard causal assumptions and the particular treatment effect heterogeneity assumptions specified. Among this class of unbiased estimators, the investigator can then identify the one with the lowest variance for an assumed working covariance structure. Analogous to the working correlation matrix used in generalized estimating equations (Zeger & Liang (1986)), this working covariance structure need not be correctly specified to maintain unbiased estimation. A misspecified structure, however, may lead to suboptimal variance and reduced power to detect an effect. This framework unifies ideas from both the observational staggered adoption methods and randomized SWT methods, providing estimators with more flexibility and interpretability under heterogeneity than SWT estimators and lower variance than staggered adoption methods.
@@
The composition of the estimator from two-by-two DID estimators serves several purposes. First, it handles unit- and time-confounding and ensures the use of both horizontal and vertical comparisons, as discussed by Matthews & Forbes (2017). Second, it allows both intuitive connections to methods commonly used in econometrics and quantitative social science for quasi-experimental panel data as well as assumptions for validity formulated similarly to those in the existing literature of that field. Third, it allows clear formulation of the treatment effect heterogeneity assumptions that are made rather than those assumptions arising implicitly through a linear regression formulation. In fact, these building blocks, including non-canonical two-by-two comparisons, have independently been proposed as useful methods to test for heterogeneity of treatment effects (Shahn & Hatfield (2024); Xu et al. (2024)). By specifying the variance assumption separately, it also removes the need for a correct joint specification of the mean and variance models and allows those assumptions (for the working structure) to be formulated independently of the treatment effect assumptions. These approaches also allow for ease of sensitivity analysis, as the expectation can be found using Equation 5 using a specified weight vector under any other assumption setting. In particular, unlike the mixed effects models that are commonly used for SWT analysis, the precise estimand targeted if there are additional unspecified heterogeneities can be identified, allowing more robust interpretation. However, this comes at a small cost to power and precision compared to a correctly-specified mixed effects model, as shown in the simulations. Along with the potential to assess relative efficiency under the working covariance structure as shown in the examples, these properties allow the bias-variance tradeoff to be made explicit, and permit investigators to consider the appropriate generalizability and interpretation of their target estimands.
@@
The example re-analyzing the data collected by Trajman et al. (2015) demonstrates the use of this method for a variety of target estimands in a SWT. Results largely align with those of existing non-parametric estimators, which already had been shown to have desirable bias and inference properties (Thompson et al. (2018); Kennedy‐Shaffer et al. (2020)). In nearly comparable situations, like the assumption of a homogeneous effect compared to the inverse-variance weighted average vertical approach (Thompson et al. (2018)) or either crossover-type estimator (Kennedy‐Shaffer et al. (2020)), this approach had a lower p-value, which may reflect improved precision, as shown in the simulations. In addition, many other estimates are tractable using this approach, allowing for flexible pre-specified or exploratory analysis of treatment effect heterogeneity.
The observational staggered adoption setting using data collected and analyzed by Fuller et al. (2022) demonstrates that this method can also be used for non-randomized settings. Results again directionally align with those identified using various synthetic control and DID estimators, but this method allows more specific targeting of certain effects while also making clear the homogeneity assumptions being made. These are largely similar to the results obtained using recently-proposed staggered adoption methods by Callaway & Sant’Anna (2021) and Sun & Abraham (2021), but allow for the use of more data by not restricting to the last-untreated period. The simulations demonstrate that, if the assumptions are met, the generalized DID estimator tends to be more efficient than corresponding estimators of the forms described by Callaway & Sant’Anna (2021) and Sun & Abraham (2021). Again, it is important to consider the parallel trends assumption here as well in order to interpret the estimates causally. Trade-offs between efficiency and validity arise due to these additional considerations in the non-randomized case, and further exploration of this specific consideration is warranted.
@@
The simulations demonstrate the feasibility of using these estimators in the analysis of stepped wedge trials and the tradeoffs compared to existing mixed effects model approaches and existing staggered adoption approaches. For non-randomized settings, simulation under a reasonable data-generating process should be conducted to confirm these general results. The improved efficiency compared to purely vertical, pure crossover approaches, or staggered adoption approaches that limit the comparisons used, however, should all remain, as this approach uses as much information as possible under the assumed treatment effect heterogeneities. The building blocks enforce zero-sums on weights in both rows and columns, as discussed by Matthews & Forbes (2017). In addition, the fact that the method generally highly weights cells with high information content (Kasza & Forbes (2019); Kasza, Taljaard & Forbes (2019)) and respects centrosymmetry under appropriate assumptions (Bowden et al. (2021)) indicates efficient use of the observed data. However, their performance compared to regression-based approaches (e.g., Hooper et al. (2016); Lindner & McConnell (2021); Kenny et al. (2022); Maleyeff et al. (2023); Lee & Cheung (2024)) and existing robust estimators (e.g., Hughes et al. (2020); Roth & Sant’Anna (2023); Borusyak et al. (2024)) remains to be determined. In particular, future study and setting-specific simulations are needed to explore when this approach provides equivalent estimators to those approaches and when one outperforms the other. Further work is also needed in the non-randomized case to determine the differences in efficiency and in the assumptions required between these approaches and existing staggered adoption estimators, extending the work in Section 4 to a variety of settings.
@@
Additional future work should consider identifying other approaches to inference, including the possibility of closed-form variance estimators and asymptotic sampling distributions under certain variance assumptions. Identifying appropriate plug-in estimators for use in Equation 11, their robustness to misspecification of the working covariance structure, and the asymptotic distribution would be key, as has been done for mixed effects models with misspecification (see, e.g., Voldal et al. (2022)). This would allow for more targeted design of experiments when analysis will proceed using these methods. Understanding likely covariance structures, especially given planned or conducted approaches for sampling of units (see, e.g., Hooper (2021)), would improve selection of $\bm{M}$ as well. This would also enable either simulation or analytic approaches to sample size determination, a key feature in study design. Additionally, modifications of this method to accommodate adjustment for covariates would be useful for non-randomized staggered adoption settings where parallel trends holds only conditionally (see, e.g., Callaway & Sant’Anna (2021); Roth et al. (2023)) and SWTs with restricted, stratified, or matched randomization (see, e.g., Copas et al. (2015)).
@@
Both stepped wedge and staggered adoption settings are common study designs with important roles to play in assessing a variety of policies and implementation approaches; for example, both have been proposed as useful tools for rapid policy evaluation in the COVID-19 pandemic or similar settings (see, e.g., Goodman-Bacon & Marcus (2020); Bell et al. (2021); Cowger et al. (2022); Kennedy-Shaffer (2024)). They each, however, have statistical intricacies that should be carefully considered in the design and analysis (see, e.g., Baker et al. (2022); Nevins et al. (2024)). The generalized DID estimator framework proposed here has desirable bias and variance properties and interpretability that enable its use across many such settings and desired treatment effect estimands.
@@
All code used for this manuscript is available at https://github.com/leekshaffer/GenDID. Included there is a simulated data set based on the stepped-wedge trial data example. Access policies for the original data set can be found in Trajman et al. (2015) at https://doi.org/10.1371/journal.pone.0123252. The original data set for the staggered adoption example are included at https://github.com/leekshaffer/GenDID; details on the collection of that data can be found in Fuller et al. (2022) at https://doi.org/10.1371/journal.pone.0274374. The original data repository for that article is available at https://doi.org/10.7910/DVN/K1XX02; the file used is synth_data_clean.Rdata. See the repository for citation, access, and sharing policies.
Acknowledgements
The author wishes to thank Prof. Anete Trajman for making the stepped-wedge trial data available to the author for re-analysis and Prof. Jennifer Thompson for helpful discussions regarding the data. The author wishes to thank Sam Fuller, Sara Kazemian, Prof. Carlos Algara, and Prof. Daniel J. Simmons for making their COVID-19 vaccine lottery data publicly available. The author also wishes to thank colleagues at Vassar College (specifically, Profs. Andy Borum, Adam Lowrance, Lisa Lowrance, and Benjamin Morin and students Raymond Aye, Ryan Kilpadi, and Andrew Steindl) for helpful discussions on linear algebra and colleagues at the Yale School of Public Health (specifically, Profs. Fan Li and Donna Spiegelman) for helpful discussions on stepped-wedge trial settings. Finally, the author wishes to thank participants in the Current Developments in Cluster Randomised Trials and Stepped Wedge Designs conference (Queen Mary University of London, November 2022), the Fourth International Conference on Stepped Wedge Design (University of York, May 2024), and the University of Illinois Department of Statistics seminar series (September 2024) for helpful feedback on this project.